Word problems always twist me around. Your assistance is greatly appreciated. Thank you in advance.
An investment club used funds totaling $300,000 to invest in a bio-tech company and in an ethanol plant, with hopes of earning 12% and 16% annual simple interest, respectively. Their hunch paid off. The club made a total of $38,600 interest the first year. How much was invested at each rate?

$____________ at 12%

$ _____________at 16%

Answer by TimothyLamb(4379)

(0.12)b + (0.16)e = 38600
b + e = 300000
b = 300000 - e
---
(0.12)b + (0.16)e = 38600
(0.12)(300000 - e) + (0.16)e = 38600
(0.12)300000 - (0.12)e + (0.16)e = 38600
(0.04)e = 2600
e = 2600/(0.04)
---
e = $65000
b = $235000
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php